OpenStudy (anonymous):

Write the equation of the line for the following condition. through two points (2, 4) and (4, 7) @shamil98

OpenStudy (shamil98):

find the slope, use y =mx + b plug in the coordinates solve for b. put your value of along with m your slope to find it

OpenStudy (anonymous):

Sorry I don't quite understand the y = mx + b formula.

OpenStudy (shamil98):

find the slope first.

OpenStudy (anonymous):

So like the same formula I used for my previous question?

OpenStudy (shamil98):

yes

OpenStudy (anonymous):

3/2

OpenStudy (shamil98):

no.. you got the x value on top and y on the bottom its the other way around

OpenStudy (anonymous):

Oh yeah my bad. 2/3

OpenStudy (anonymous):

Wait,

OpenStudy (anonymous):

Yeah 2/3

OpenStudy (shamil98):

yes

OpenStudy (shamil98):

input your values of (x,y) and m into y=mx + b what do u get,

OpenStudy (anonymous):

2 = m3 + b ...? Sorry I'm not so good at this.

OpenStudy (shamil98):

nvm, i got the slope wrong, its 3/2

OpenStudy (anonymous):

So I had it right before?

OpenStudy (shamil98):

(x,y) --> (2,4) m = 3/2 y = mx + b 4 = (3/2)2 + b yes you did

OpenStudy (shamil98):

solve for b.

OpenStudy (anonymous):

b = 1?

OpenStudy (anonymous):

Because I got 4 = 3 + b

OpenStudy (shamil98):

yes

OpenStudy (shamil98):

y = mx +b y = 3/2 x + 1

OpenStudy (anonymous):

so 'm' is the slope right? For the formula y=mx + b

OpenStudy (shamil98):

yeah

OpenStudy (anonymous):

Okay now I understand the formula... But there's another question, can you help me with it? I'll put it on another thread.

OpenStudy (shamil98):

ok